# Measuring Visual Sycophancy in Multimodal Models

### v1

@@ 5 Related Work
Our study on visual sycophancy in multimodal language models builds upon several key research areas. Recent advancements in multimodal AI systems have expanded the capabilities of models to process and integrate information from various modalities (Wu et al., 2024; Ge et al., 2024; Li et al., 2024a, b). This progress has been accompanied by growing concerns about biases in AI, including both social and modality-specific biases (Luo et al., 2024; Alabdulmohsin et al., 2024; Lu et al., 2024; Adewumi et al., 2024; Chen et al., 2024).
Visual attention mechanisms in AI systems have been studied extensively, often drawing parallels with human visual processing (Cao et al., 2024). Evaluation methodologies for multimodal AI systems have evolved to address the complexities of assessing performance across different modalities (Ye et al., 2024). Simultaneously, ethical considerations in AI development have gained prominence, focusing on the potential impacts of AI biases in real-world applications (Amirloo et al., 2024).
@@ 6 Limitations and Future Work
Although we investigate and analyze a few of the most well-known proprietary and open-source models, the generalizability of our findings could be enhanced by including a broader range of state-of-the-art models. Future work should aim to test these concepts across an even more diverse set of architectures and configurations. Since multimodal prompt engineering and modality bias are relatively new concepts, our primary focus has been on measuring the bias rather than proposing specific applications or effective jailbreak mitigation strategies. Further research is needed to develop practical interventions and assess their effectiveness in real-world scenarios.
The use of token probability delta as a novel metric for calculating bias in machine learning models is still in its early stages. It is not yet entirely clear whether systematic bias in multimodal machine learning models is inherently additive or subtractive, and this remains an area for further empirical and theoretical investigation. Modalities other than vision and text have not been explored in this study. Future research should consider extending the analysis to include other modalities, such as audio or sensor data, to determine whether visual sycophancy or similar biases are present across different types of multimodal inputs.
Our experiments primarily focus on visual biases within the context of multiple-choice benchmarks. The extent to which these findings translate to more complex, free-text generation tasks or other forms of human-AI interaction remains unexplored and could be an avenue for future research. While our study addresses the phenomenon of visual sycophancy, we have not fully explored the potential interactions between visual biases and other types of biases (e.g., social or cognitive biases) present in multimodal models. Understanding these interactions could provide a more comprehensive picture of bias in AI systems.
@@ 7 Conclusion
Our study on visual sycophancy in multimodal language models reveals a complex landscape of model behaviors and biases. We found that the susceptibility to visual cues varies significantly across different model architectures, task types, and visual presentation formats.
Our findings challenge simplistic assumptions about multimodal information integration in AI systems and raise important questions about the reliability and consistency of model outputs. The observed visual sycophancy effects underscore the need for careful consideration of visual elements in the design and deployment of multimodal AI systems, particularly in critical decision-making contexts.

### v2

@@ 5 Related Work
Our study on agreeableness bias in multimodal language models builds upon several key research areas. Recent advancements in multimodal AI systems have expanded the capabilities of models to process and integrate information from various modalities (Wu et al., 2024; Ge et al., 2024; Li et al., 2024a, b). This progress has been accompanied by growing concerns about biases in AI, including both social and modality-specific biases (Luo et al., 2024; Alabdulmohsin et al., 2024; Lu et al., 2024; Adewumi et al., 2024; Chen et al., 2024). Visual attention mechanisms in AI systems have been studied extensively, often drawing parallels with human visual processing (Cao et al., 2024). Evaluation methodologies for multimodal AI systems have evolved to address the complexities of assessing performance across different modalities (Ye et al., 2024). Simultaneously, ethical considerations in AI development have gained prominence, focusing on the potential impacts of AI biases in real-world applications (Amirloo et al., 2024).
@@ 6 Limitations and Future Work
Although we investigate and analyze a few of the most well-known proprietary and open-source models, the generalizability of our findings could be enhanced by including a broader range of state-of-the-art models. Future work should aim to test these concepts across an even more diverse set of architectures and configurations. Since multimodal prompt engineering and agreeableness bias are relatively new concepts, our primary focus has been on measuring the bias rather than proposing specific applications or effective jailbreak mitigation strategies. Further research is needed to develop practical interventions and assess their effectiveness in real-world scenarios.
The use of token probability delta as a novel metric for calculating bias in machine learning models is still in its early stages. It is not yet entirely clear whether systematic bias in multimodal machine learning models is inherently additive or subtractive, and this remains an area for further empirical and theoretical investigation. Modalities other than vision and text have not been explored in this study. Future research should consider extending the analysis to include other modalities, such as audio or sensor data, to determine whether agreeableness bias or similar biases are present across different types of multimodal inputs.
Our experiments primarily focus on visual biases within the context of multiple-choice benchmarks. The extent to which these findings translate to more complex, free-text generation tasks or other forms of human-AI interaction remains unexplored and could be an avenue for future research. While our study addresses the phenomenon of agreeableness bias, we have not fully explored the potential interactions between visual biases and other types of biases (e.g., social or cognitive biases) present in multimodal models. Understanding these interactions could provide a more comprehensive picture of bias in AI systems.
@@ 7 Conclusion
Our study on agreeableness bias in multimodal language models reveals a complex landscape of model behaviors and biases. We found that the susceptibility to visual cues varies significantly across different model architectures, task types, and visual presentation formats. Our findings challenge simplistic assumptions about multimodal information integration in AI systems and raise important questions about the reliability and consistency of model outputs. The observed agreeableness bias effects underscore the need for careful consideration of visual elements in the design and deployment of multimodal AI systems, particularly in critical decision-making contexts.